The solution of ye^{x}dx + (2y + e^{x})dy=0, y(0) = -1 is A) y = e^{x}. B) y=e^{-x} C) y = 1 D) y = -1


1
Expert's answer
2021-08-20T01:37:06-0400

"ye^xdx+\left(2y+e^x\right)dy=0\\
\mathrm{An\:ODE\:}M\left(x,\:y\right)+N\left(x,\:y\right)y'=0\mathrm{\:is\:in\:exact\:form\:if\:the\:following\:holds:}\\
1.\:\:\mathrm{There\:exists\:a\:function\:}\Psi \left(x,\:y\right)\mathrm{\:such\:that\:}\Psi _x\left(x,\:y\right)=M\left(x,\:y\right),\:\quad \Psi _y\left(x,\:y\right)=N\left(x,\:y\right)\\
2.\:\:\Psi \left(x,\:y\right)\mathrm{\:has\:continuous\:partial\:derivatives:\quad }\frac{\partial M\left(x,\:y\right)}{\partial y}=\frac{\partial ^2\Psi \left(x,\:y\right)}{\partial y\partial x}=\frac{\partial ^2\Psi \left(x,\:y\right)}{\partial x\partial y}=\frac{\partial N\left(x,\:y\right)}{\partial x}\\
\mathrm{Let\:}y\mathrm{\:be\:the\:dependent\:variable.\:Divide\:by\:}dx\mathrm{:}\\
ye^x+\left(2y+e^x\right)\frac{dy}{dx}=0\\
\mathrm{Substitute\quad }\frac{dy}{dx}\mathrm{\:with\:}y'\:\\
ye^x+\left(2y+e^x\right)y'\:=0\\
\mathrm{If\:the\:conditions\:are\:met,\:then\:}\Psi _x+\Psi _y\cdot \:y'=\frac{d\Psi \left(x,\:y\right)}{dx}=0\\
\mathrm{The\:general\:solution\:is\:}\Psi \left(x,\:y\right)=C\\
\mathrm{True}\\
Ψ\left(x,\:y\right)=c_2\\
e^xy+y^2+c_1=c_2\\
\mathrm{Combine\:the\:constants}\\
e^xy+y^2=c_1\\
y=\frac{-e^x+\sqrt{e^{2x}+4c_1}}{2},\:y=\frac{-e^x-\sqrt{e^{2x}+4c_1}}{2}\\
y=\frac{-e^x+\sqrt{e^{2x}+c_1}}{2},\:y=\frac{-e^x-\sqrt{e^{2x}+c_1}}{2}"